At a depth of 1000m in an ocean (a)what is the absolute pressure?(b)what is the gauge pressure?PLEASE ANSWER FAST.....

Grade:11

1 Answers

Arun
6 years ago

(a) Absolute Pressure = Padgh = 105 + (103)(10)(1000) = 10.1 x 10N/m2

(b) Gauge Pressure = (103)(10)(1000) = 10 x 10N/m2